# On Bits and Bandits: Quantifying the Regret-Information Trade-off

### v2

@@ Appendix D Proofs for Section 4
We begin by defining the W-Lambert function and supplying useful properties of it.
Definition D.1.
The Lambert W-function is $W:[-1,\infty]\to[-1,\infty]$ is the inverse function to $x2^{x}$. This means that for all $x\geq 0$, $y=x2^{x}$ if and only if $W(y)=x$.
We now state the following useful property of the W-Lambert function.
Lemma D.1 (Equation 21 of [Iacono and Boyd, 2017]).
$W(x)=\Theta\left(\log\left(\frac{x}{\log x}\right)\right)$
We begin by proving the following useful lemma.
@@
Assume a finite decision space with $n$ decisions. Then the highest $L_{1}$ norm of every decision in the set, under the mutual information constraint of $I(\pi^{*};\mathcal{H}_{T})\leq R\leq H(\pi^{*})$ is at most $\Theta\left(\frac{R}{\log n}\right)$.
@@
We begin by noting that the mutual information can be written as $I(\pi^{*};\mathcal{H}_{T})=D\left(P^{\pi^{*}\mid\mathcal{H}_{T}},P^{\pi^{*}}\right)$. Thus, the constrained decision space becomes the decisions space under the KL constraint, which can be written as
$\sum_{i}\phi(i)\log\frac{\phi(i)}{P(i)}\leq R.$
Thus, we have that
$\phi^{*}\log\frac{\phi^{*}}{p}\leq R$
we re-parameterize $\psi=\frac{\phi^{*}}{p}$ and obtain
$p\psi\log\psi\leq R.$
Hence,
$\psi\leq 2^{W(R/p)}=\Theta\left(\frac{R}{p\log\left(R/p\right)}\right).$
We substitute $\phi^{*}$ into the expression,
$\phi^{*}\leq\Theta\left(\frac{R}{\log\left(R/p\right)}\right)=\Theta\left(\frac{R}{\log n}\right).$
We now state the following lemma that explains how we can use the results of Lemma D.2 to Equation 4.
Lemma D.3.
Under the assumptions and notations of Proposition 3.2, and under the constraint of $I(\pi^{*};\mathcal{H}_{T})\leq R\leq H(\pi^{*})$, the following holds.
$I(V;\mathcal{H}_{T})\lesssim 2T\frac{R}{\log K}\epsilon^{2}.$
Proof.
The proof follows a similar analysis to the proof of Theorem 3.4, which is provided in Yang and Barron [1999]. We denote $E$ as the space of the constrained distributions. We now have
$I(V;\mathcal{H}_{T})$ $\leq T\max_{p_{1},p_{2}\in E}D\left(p_{1},p_{2}\right)$
$\lesssim T\frac{R}{\log K}\max_{p_{1},p_{2}\in E}\left\|p_{1}-p_{2}\right\|_{2}^{2}$
$\leq T\frac{R}{\log K}\epsilon^{2}.$
where the second inequality follows from D.2. ∎
@@
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\gtrsim 1-\frac{2\epsilon^{2}T\frac{R}{\log K}+1}{\log\mathcal{M}_{1}(\Delta(\Pi),\epsilon/2)-\log\mathcal{M}_{1}(\Delta(\Pi),\epsilon)}$
@@
The proof follows immediately by utilizing Lemma D.3 and using the fact that in this scenario, $\rho(x,y)=\left\|x-y\right\|_{1}$, similarly to the proof in Lemma C.1. ∎
@@
We begin by denoting the polytope space given the mutual information constraint as $\Delta_{R}$. We first see that
$\log\mathcal{M}_{1}(\epsilon/2,\Delta_{R})-\log\mathcal{M}_{1}(\epsilon,\Delta_{R})=\Theta\left(K\right).$
@@
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\gtrsim 1-\frac{2T\frac{R}{\log K}\epsilon^{2}+1}{K}.$
Selecting $\epsilon=\frac{1}{4}\sqrt{\frac{RK}{T\log K}}$ we now have that
$\mathcal{BR}^{*}(T)\gtrsim\frac{1}{2}\sqrt{TK\frac{R}{\log K}}$
@@
$\log\mathcal{M}_{1}(\epsilon/2,\Delta_{R})-\log\mathcal{M}(\epsilon,\Delta_{R})=\Theta\left(d\right).$
@@
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\gtrsim 1-\frac{2T\frac{R}{\log K}\epsilon^{2}+1}{d}$
selecting $\epsilon=\frac{1}{4}\sqrt{\frac{Rd}{T\log K}}$ we now have that
$\mathcal{BR}^{*}(T)\gtrsim\frac{1}{2}\sqrt{\frac{dTR}{\log K}}.$
@@
For a $K$-Bayesian MAB, the regret can be lower bounded by $\Omega\left(\sqrt{KT\frac{I(\pi^{*};\mathcal{H}_{T})}{\log K}}\right)$. If the MAB is also linear, then the lower bound becomes $\Omega\left(\sqrt{dT\frac{I(\pi^{*};\mathcal{H}_{T})}{\log K}}\right)$. For a Contextual MAB with $C$ contexts, the lower bound becomes $\Omega\left(\sqrt{KCT\frac{I(\pi^{*};\mathcal{H}_{T})}{\log K}}\right)$.
Appendix E Experiments Details
Figure 5: The prompting done for the LLMs in our experiments. The question is preceded by "Answer the following question:". Following this, the options are presented and to improve accuracy, the LLM is requested to output only the letter for the correct answer. Then, for the next round’s prompt, we add the LLM’s answer and feedback.
E.1 Bayesian Bandit Experiments
@@
E.2 MCQA with LLM Experiment
@@
For the second experiment, we used two different LLMs. The first is Mistral 7B Jiang et al. [2023], an LLM with 7.3B parameters. The second is Mixtral 8x7B Jiang et al. [2024]. The architecture of this model is the same as Mistral 7B, but it uses 8 of them and a mixture of experts approach to select which one to use. For both models, we use 4bit quantization of type nf4 Dettmers and Zettlemoyer [2023]. In every round, we queried the LLM with all previous prompts along with a prompt that consists of the following structure: "Answer the following question: [question]? These are the possible answers [options]. Please output only the letter for the correct answer." Using the LLM’s output we obtain a probability distribution over the tokens of the possible answers [A,B,C,D]. We then selected the answer according to this distribution and received feedback as to whether the answer was correct or not. The feedback was also added to the prompt. The entire procedure is described in Figure 5.
We repeated this process for 10 different question seeds. For each question seed, a different set of 200 questions was selected. Both policies were evaluated on each set for 10 different seeds, so 100 random seeds were used for this experiment. Figure 4 describes the behavior of both policies across 10 different seeds but using the same question seed. Table 2 reports the results across all 100 seeds.
E.3 Compute Specifications

### v3

@@ Appendix D Proofs for Section 4
We begin by stating and proving the following Lemma.
@@
Under the assumptions and notations of Proposition 3.2, the constraint of $I(\pi^{*};\mathcal{H}_{T})\leq R\leq\log K$, and the regularity Assumption 3.1, the following holds.
$I(V;\mathcal{H}_{T})\leq 2\bar{A}T\frac{R}{\log K}\epsilon^{2}.$
@@
The proof follows a similar analysis to the proof of Theorem 3.3, which is provided in Yang and Barron [1999]. Since we focus on a parametric set, we consider a local packing set with a cardinality of $\mathcal{M}^{\mathrm{loc}}(\rho,\epsilon,\Delta(\Pi))$, which we denote by $E$. $\tilde{E}$ is the set contained by $E$ under the mutual information constraint,
$\tilde{E}=\{p\in E:I(\pi^{*};\mathcal{H}_{T})\leq R\}.$
We now have
$I(V;\mathcal{H}_{T})$ $=\sum_{t=1}^{T}I(V;C_{t+1},O_{t+1},\dots,C_{T},O_{T}\mid\mathcal{H}_{t})$
$=\sum_{t=1}^{T}D\left(P^{V\mid\mathcal{H}_{T}},P^{V\mid\mathcal{H}_{t}}\right)$
$\leq T\max_{P_{1},P_{2}\in\tilde{E}}D\left(P_{1},P_{2}\right)$
$\leq T\frac{R}{\log K}\max_{P_{1},P_{2}\in E}D\left(P_{1},P_{2}\right)$
$\leq T\frac{R}{\log K}\bar{A}\max_{P_{1},P_{2}\in E}\rho(P_{1},P_{2})^{2}$
$\leq T\frac{R}{\log K}\bar{A}\epsilon^{2}.$
where the second inequality follows from a scaling argument, and the last follows from the fact that $E$ is a local packing set. ∎
@@
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\geq\frac{1}{2}\left[1-\frac{2\bar{A}\epsilon^{2}T\frac{R}{\log K}+1}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)}\right]$
Furthermore, if the decision-making problem is a part of a parametric set, then
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\geq\frac{1}{2}\left[1-\frac{2\bar{A}\epsilon^{2}T\frac{R}{\log K}+1}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Pi,\epsilon)}\right]$
@@
The proof follows immediately by utilizing Lemma D.1 and using the fact that in this scenario, $\rho(x,y)=\left\|x-y\right\|_{1}$, similarly to the proof of Proposition C.1. ∎
@@
We start by considering $R=0$. From Lemma 4.2,
$\mathcal{BR}^{*}(T)\geq\frac{\epsilon T}{2}\left(1-\frac{1}{K}\right)$
for any $\epsilon>0$ such that lower bounds the difference between two different decisions. Since the maximal difference between two decisions is $1$, we set $\epsilon=1$ and obtain the linear lower bound,
$\mathcal{BR}^{*}(T)\geq\frac{T}{2}\left(1-\frac{1}{K}\right).$
We now consider $R>0$. We denote the polytope space with the mutual information constraint by $\Delta_{R}$. Similarly to the proof of Proposition C.1,
$\log\mathcal{M}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)\geq(K-1).$
@@
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\geq\frac{1}{2}\left[1-\frac{2T\bar{A}\frac{R}{\log K}\epsilon^{2}+1}{(K-1)}\right].$
Selecting $\epsilon=\sqrt{\frac{(K-1)\log K}{6RT\bar{A}}}$ we now have that
$\mathcal{BR}^{*}(T)$ $\geq\frac{1}{2}\sqrt{\frac{TK\log K}{6R\bar{A}}}\left(\frac{2}{3}-\frac{1}{K-1}\right)$
@@
$\log\mathcal{M}^{\mathrm{loc}}(\rho,\Pi,\epsilon)\geq d.$
@@
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\geq\frac{1}{2}\left[1-\frac{2T\frac{R}{\log K}\bar{A}\epsilon^{2}+1}{d}\right].$
Selecting $\epsilon=\sqrt{\frac{d\log K}{6R\bar{A}T}}$ we now have that
$\mathcal{BR}^{*}(T)$ $\geq\frac{1}{2}\sqrt{\frac{Td\log K}{6R\bar{A}}}\left(\frac{2}{3}-\frac{1}{d}\right)$
Proposition 4.1 can then be applied directly on the MAB setting.
@@
Consider a Bayesian MAB with $K$ arms. If $I(\pi^{*};\mathcal{H}_{T})=0$ then the worst-case regret is linear. Otherwise, for $I(\pi^{*};\mathcal{H}_{T})\geq 0$, the worst-case regret can be lower-bounded by $\Omega\left(\sqrt{\frac{TK\log K}{I(\pi^{*};\mathcal{H}_{T})}}\right)$. Additionally, if $\Pi=\mathbb{B}_{2}^{d}$ and the reward means are linear functions of the decisions, we have a lower bound of $\Omega\left(\sqrt{\frac{Td\log K}{I(\pi^{*};\mathcal{H}_{T})}}\right)$.
D.2 Proof of Proposition 4.3
Proof.
We consider the following two algorithms. The first is Thompson sampling, whose policy at round $t$ is the posterior probability, $p_{t}=P^{\pi^{*}\mid\mathcal{H_{t}}}$. The second algorithm is Thompson sampling with the constrained observations $\tilde{O}_{t}$ such that $I(\pi^{*};\tilde{\mathcal{H}}_{T})=D\left(P^{\pi^{*}\mid\tilde{\mathcal{H}}_{T}},P^{\pi^{*}}\right)=R$, where ~$\tilde{\mathcal{H}}_{t}$ is the history with constrained observations. Now, we define the information ratio for both the normal and constrained variants of Thompson sampling:
$\rho_{t}=\frac{\mathbb{E}[r_{M}(\pi^{*}(C_{t}),C_{t})-r_{M}(p_{t},C_{t})]^{2}}{I(\pi^{*};O_{t+1}\mid\mathcal{H}_{t})}\quad;\quad\tilde{\rho}_{t}=\frac{\mathbb{E}[r_{M}(\pi^{*}(C_{t}),C_{t})-r_{M}(p_{t},C_{t})]^{2}}{I(\pi^{*};\tilde{O}_{t+1}\mid\tilde{\mathcal{H}_{t}})}.$
We see that
$\frac{\rho_{t}}{\tilde{\rho}_{t}}$ $=\frac{I(\pi^{*};\tilde{O}_{t+1}\mid\tilde{\mathcal{H}_{t}})}{I(\pi^{*};O_{t+1}\mid\mathcal{H}_{t})}$
$\geq\frac{I(\pi^{*};\tilde{O}_{t+1}\mid\tilde{\mathcal{H}_{t}})}{\log K}$
where the inequality follows from $I(\pi^{*};O_{t+1}\mid\mathcal{H}_{t})\leq\log K$. Thus,
$\max_{P}\max_{t}\tilde{\rho}_{t}$ $\leq\max_{P}\max_{t}\frac{\log K}{D\left(P^{\pi^{*}\mid(\tilde{O_{t}},\tilde{\mathcal{H}}_{t})},P^{\pi^{*}\mid\tilde{\mathcal{H}}_{t}}\right)}\rho_{t}$
$\leq\frac{\log K}{R}\max_{t}\rho_{t}$
$\leq\frac{K\log K}{2R}$
Where the third inequality follows by the information ratio upper bound presented by Russo and Van Roy [2016]. If the MAB is linear then the information ratio $\rho_{t}$ can be upper-bounded by $\frac{d}{2}$. Then, we have the following upper bound,
$\max_{P}\max_{t}\tilde{\rho}_{t}\leq\frac{d\log K}{2R}$
Now, by the analysis done by Russo and Van Roy [2016], we know that if $\max_{t}\tilde{\rho}_{t}\leq\tilde{\rho}$ for all $t$, then the Bayesian regret can be upper bounded by
$\mathcal{BR}^{*}(T)$ $\leq\sqrt{\tilde{\rho}T\log K}$
Thus, we have for $K$-MAB:
$\mathcal{BR}^{*}(T)\leq\log K\sqrt{\frac{KT}{2R}}.$
And for linear bandits:
$\mathcal{BR}^{*}(T)\leq\log K\sqrt{\frac{dT}{2R}}.$
D.3 Proof of Proposition 4.5
We begin by stating and proving the following Lemma.
Lemma D.4.
Let $P_{1},P_{2}$ be two priors such that $H(P_{1}^{\pi^{*}})=R\leq\log K$. Also let $\psi,\phi$ be two decisions such that $D\left(\psi,P_{1}^{\pi^{*}}\right)\geq c\bar{A}R$ and $D\left(\phi,P_{2}^{\pi^{*}}\right)\leq\log K$. Then,
$\frac{\mathbb{E}_{P_{1}}[\mu_{M}(\pi^{*},C)-\mu_{M}(\psi,C)]}{\mathbb{E}_{P_{2}}[\mu_{M}(\pi^{*},C)-\mu_{M}(\phi,C)]}\geq 2\sqrt{c\frac{R}{\log K}}$
Proof.
By the regularity assumption, we know that
$(\mathbb{E}_{P_{1}}[\mu_{M}(\pi^{*},C)-\mu_{M}(\psi,C)])^{2}\geq\frac{cR}{2}.$
From Pinsker’s inequality,
$(\mathbb{E}_{P_{2}}[\mu_{M}(\pi^{*},C)-\mu_{M}(\phi,C)])^{2}\leq 2D\left(\phi,P_{2}^{\pi^{*}}\right)\leq\log K.$
Dividing these two inequalities yields
$\frac{(\mathbb{E}_{P_{1}}[\mu_{M}(\pi^{*},C)-\mu_{M}(\psi,C)])^{2}}{(\mathbb{E}_{P_{2}}[\mu_{M}(\pi^{*},C)-\mu_{M}(\phi,C)])^{2}}\geq 4c\frac{R}{\log K}$
which concludes our proof. ∎
Now, we prove the following Proposition.
Proposition D.5.
Let $P$ be a prior $P$ with $H(\pi^{*})=R\leq\log K$. If Assumption 4.1 holds, then
$\mathcal{BR}_{P}(T;\{p_{t}\})\geq 2\sqrt{\frac{cR}{\log K}}\mathcal{BR}^{*}(T).$ (15)
Proof.
We see that
$\mathcal{BR}_{P}(T;{p_{t}})$ $=\sum_{t=1}^{T}\mathbb{E}_{P}[\mu_{M}(\pi^{*},C_{t})-\mu_{M}(p_{t},C_{t})]$
$\geq 2\sqrt{c\frac{R}{\log K}}\sum_{t=1}^{T}\mathbb{E}_{\nu}[\mu_{M}(\pi^{*},C_{t})-\mu_{M}(p_{t}^{\prime},C_{t})]$
for any prior $\nu$ and policies $p^{\prime}_{t}$ such that $D\left(p^{\prime}_{t},\nu^{\pi^{*}}\right)\leq\log K$. So we can take the worst-case regret,
$\mathcal{BR}_{P}(T;{p_{t}})$ $\geq 2\sqrt{c\frac{R}{\log K}}\max_{\nu}\inf_{\{p^{\prime}_{t}\}:D\left(p^{\prime}_{t},\nu^{\pi^{*}}\right)\leq\log K}\sum_{t=1}^{T}\mathbb{E}_{\nu}[\mu_{M}(\pi^{*},C_{t})-\mu_{M}(p_{t}^{\prime},C_{t})]$
$\geq 2\sqrt{c\frac{R}{\log K}}\max_{\nu}\inf_{\{p^{\prime}_{t}\}}\sum_{t=1}^{T}\mathbb{E}_{\nu}[\mu_{M}(\pi^{*},C_{t})-\mu_{M}(p_{t}^{\prime},C_{t})]$
$=2\sqrt{c\frac{R}{\log K}}\mathcal{BR}^{*}(T)$
where the second inequality follows by removing a constraint on the available policies. ∎
This results in the following Corollary.
Corollary D.6 (Proposition 4.5).
Let there be a $K$-MAB. Then for any algorithm such that $D\left(p_{t},P^{\pi^{*}}\right)\geq c\bar{A}H(\pi^{*})$, the regret can be lower-bounded by $\Omega\left(\sqrt{\frac{KH(\pi^{*})T}{\log K}}\right)$. Additionally, if the MAB is linear, then the regret can be lower bounded by $\Omega\left(\sqrt{\frac{dH(\pi^{*})T}{\log K}}\right)$.
We now explain why in scenarios where the optimal decision is unique, an algorithm that selects the optimal arm with high probability obeys Assumption 4.1.
Proposition D.7.
Let there be an algorithm such that $p_{t}$ selects the optimal decision $\pi^{*}$ with probability $1-\delta$. Then,
$D\left(p_{t},P^{\pi^{*}}\right)\geq(1-\delta)H(\pi^{*})-c_{\delta}$
where $c_{\delta}\to 0$ as $\delta\to 0$.
Proof.
We can write that
$D\left(p_{t},P^{\pi^{*}}\right)$ $=\mathbb{E}\left[\sum_{\pi}p_{t}(\pi)\log\left(\frac{p_{t}(\pi)}{P^{\pi^{*}}(\pi)}\right)\right]$
$\geq\mathbb{E}_{P}\left[(1-\delta)\log\left(\frac{1-\delta}{P^{\pi^{*}}(\pi^{*})}\right)\right]$
$=(1-\delta)H(\pi^{*})-(1-\delta)\log\left(\frac{1}{1-\delta}\right)$
where the second inequality follows from the fact that $p_{t}(\pi^{*})\geq 1-\delta$. ∎
From this, we can conclude that any algorithm whose policy converges to the optimal decision obeys the assumption.
(a) Bayesian regret
(b) Accumulated information
Appendix E Bandit Experiments
E.1 Additional Details
@@
E.2 Additional Experiments
We use the same bandit setting as in Section 5. In this experiment, we compare the regret of Thompson sampling under three $H(\pi^{*})$ values. We use the same model space and number of arms. Results are presented in Figure 3. As we can see, when less information can be accumulated, less regret is suffered.
Figure 4: The prompting done for the LLMs in our experiments. The question is preceded by a question prompt followed by "Answer the following question:". Following this, the options are presented. <Question prompt> and <End prompt> are both replaced by a different prompt for every model.
Appendix F MCQA with LLM Experiment
F.1 Additional details
@@
In this experiment, we used a small LLM and a large LLM in every experiment. The small LLM was one of the following: Mistral 7B [Jiang et al., 2023], Gemma 2B [Team et al., 2024] and Falcon 7B [Almazrouei et al., 2023]. For the large LLM we used Mixtral 8x7B [Jiang et al., 2024], Llama3-70B [AI@Meta, 2024] and Gemma 7B [Team et al., 2024]. 4-bit quantization [Dettmers and Zettlemoyer, 2023] and flash attention [Dao et al., 2022] were applied for all models (excluding Falcon 7B, for which flash attention was not applied). In every round, we queried the LLM with a prompt, which is described later on. An example of the prompt used is provided in Figure 4. Using the LLM’s output we obtain a probability distribution over the tokens of the possible answers [A,B,C,D]. We then selected the answer according to this distribution and received feedback on whether the answer was correct, which was our reward. We also see that performance improvement is not clear when we use Mistral for the small model and Gemma 7B for the large one. This is because they are similar in size, while for the rest of the experiments, the larger LLM has significantly more parameters.
Table 4: Regret for 10 different seeds for random and bits-based policies in LLM selection for MCQA tasks. Both policies were allowed to query the large LLM for only 50% of the episodes. The top row corresponds with the large LLM and the left column with the small one.
Random Bits-Based
Mixtral Llama3 Gemma 7B Mixtral Llama3 Gemma 7B
Mistral $38.7\pm 0.3$ ${47.0\pm 0.5}$ $40.4\pm 0.3$ $\mathbf{36\pm 1}$ $\mathbf{46\pm 1}$ $\mathbf{40\pm 1}$
Gemma 2B $44.5\pm 0.7$ $54.1\pm 0.6$ $39.3\pm 0.3$ $\mathbf{38\pm 2}$ $\mathbf{46\pm 2}$ $\mathbf{35.3\pm 0.9}$
Falcon 7B $44.0\pm 0.3$ $52.2\pm 0.4$ $45.5\pm 0.3$ $\mathbf{35.0\pm 0.8}$ $\mathbf{44\pm 1}$ $\mathbf{39\pm 1}$
Table 5: Regret for 10 different seeds for random and bits-based policies in LLM selection for MCQA tasks. Both policies were allowed to query the large LLM for only 70% of the episodes. The top row corresponds with the large LLM and the left column with the small one.
Random Bits-Based
Mixtral Llama3 Gemma 7B Mixtral Llama3 Gemma 7B
Mistral $30.5\pm 0.4$ ${33.8\pm 0.4}$ $34.8\pm 0.3$ $\mathbf{26.6\pm 0.8}$ $\mathbf{31\pm 1}$ $\mathbf{33.0\pm 0.8}$
Gemma 2B $35.6\pm 0.8$ $40.1\pm 0.6$ $32.9\pm 0.3$ $\mathbf{32\pm 1}$ $\mathbf{36\pm 2}$ $\mathbf{30\pm 1}$
Falcon 7B $34.0\pm 0.3$ $37.0\pm 0.3$ ${37.9\pm 0.2}$ $\mathbf{27.0\pm 0.7}$ $\mathbf{31\pm 1}$ $\mathbf{33\pm 1}$
We repeated this process for 10 different question seeds. Table 2 reports the regret after 200 steps for all combinations of LLMs across all 10 seeds. Additional results are presented in Tables 4 and 5. Both tables compare the random selection method and the bits-based one for different LLMs. Table 4 compares the regret for 50% deployment rate of the large LLM and Table 5 compares the regret for 70% deployment rate of the large LLM. We conclude from both tables that the effect of the bits-based policy increases with the deployment rate of the large LLM. Furthermore, we also see that bits-based selection becomes more significant, as the performance gap between the small and large LLM increases as well.
F.2 Compute Specifications